OpenStudy (anonymous):

If y = (x + 3)2, and the value of x=5, what is the value?

OpenStudy (anonymous):

ok Alley we will break this down. are you ready?

OpenStudy (anonymous):

If Tom buys a red skateboard, for $50 and then Anne buys the same skateboard for 25% off. How much did Anne pay for her skateboard?

OpenStudy (anonymous):

If y = (x + 3)2, and the value of x=5, what is the value? y=(5+3)2 y=10+6 y=16ans.

OpenStudy (anonymous):

If Tom buys a red skateboard, for $50 and then Anne buys the same skateboard for 25% off. How much did Anne pay for her skateboard? Therefore; 1 Skateboard cost =$50 Anne got hers for 25% off = $50x25/100=$12.5 then Anne's got the skateboard for = $50-$12.5= $37.5ans.

OpenStudy (anonymous):

Deb and Sam go to the mall, each have $200 to spend. Deb spent $150 and Sam spent $160. What percent of their total money did they spend? A. 57.5% B. Cannot calculate the percentage with the information provided. C. 77.5% D. 75.7%

OpenStudy (anonymous):

Deb and Sam go to the mall, each have $200 to spend. Deb spent $150 and Sam spent $160. What percent of their total money did they spend? For Deb; $150/$200 x 100= 75% For Sam; $160/$200 x 100= 80% , so therefore the options does not contain the the answer.